# Improved SINR Approximation for Downlink SDMA-based Networks with Outdated Channel State Information

### v1

@@ Abstract
Index Terms:
Massive MIMO, rate-splitting, outdated CSI, sum-rate, performance.
00footnotetext: Manuscript received August 13, 2025; revised XXX. This work was supported in part by grant PID2023-149975OB-I00 (COSTUME) funded by MICIU/AEI/10.13039/501100011033 and FEDER/UE, in part by grant DGP_PRED_2024_0212 funded by Consejería de Universidad, Investigación e Innovación of Junta de Andalucía and the European Union, and in part by grant PMA1-2025-14-017 funded by Asociación Universitaria Iberoamericana de Postgrado. The review of this paper was coordinated by $XX$. (Corresponding author: M.C. Fernández Montefiore)00footnotetext: This work has been submitted to the IEEE for publication. Copyright may be transferred without notice, after which this version may no longer be accessible.00footnotetext: M.C. Fernández Montefiore and F.J. López-Martínez are with the Dept. Signal Theory, Networking and Communications, Research Centre for Information and Communication Technologies (CITIC-UGR), University of Granada, 18071, Granada, Spain. M.C. Fernández Montefiore, G. González and F. Gregorio are with Instituto de Investigaciones en Ingeniería Eléctrica “Alfredo Desages” (IIIE), UNS-CONICET, and Departamento de Ingeniería Eléctrica y de Computadoras, Universidad Nacional del Sur (UNS), Argentina.00footnotetext: Digital Object Identifier 10.1109/XXX.2021.XXXXXXX
@@ I Introduction
In MU-MIMO systems, rate-splitting multiple access (RSMA) is proposed as an effective strategy to overcome the challenges posed by imperfect CSIT [3]. In contrast to conventional multiple access techniques like space-division multiple access (SDMA), which treats interference as noise, and non-orthogonal multiple access (NOMA), leaning on interference decoding and cancellation, RSMA splits user messages into common and private parts, providing a more flexible approach [6] and a seamless integration with multiantenna operation. While private parts are dealt individually, the common message is decoded by all users. This allows RSMA to achieve a balance between interference suppression and decoding [7]. Furthermore, depending on the network topology and channel state condition, RSMA can adjust the power and message allocation based on interference levels, naturally transitioning between SDMA and NOMA [1, 8]. Such flexibility in managing interference makes RSMA more robust than conventional approaches, enabling a better mitigation of multi-user interference [9, 10] even in the presence of imperfect CSIT.
@@
Notation: $f_{(\cdot)}(\cdot)$ denotes a probability density function (PDF); $a$ is a scalar, $\mathbf{a}$ is a vector, and $\mathbf{A}$ is a matrix. Conjugate transpose of $\mathbf{A}$ is denoted by $\mathbf{A}^{\rm{H}}$. $\overline{a}$ denotes the complex conjugate of $a$. The expectation operator is denoted as $\mathbb{E}\left\{\cdot\right\}$, and $\mathrm{Var}\{\cdot\}$ denotes variance. $\mathcal{C}\mathcal{N}(\mu,\sigma^{2})$ represents the complex Gaussian distribution with parameters $\mu$ and $\sigma^{2}$; $\mathcal{G}(D,\Theta)$ represents the Gamma distribution with scale and shape parameters $D$ and $\Theta$, respectively. The symbol $\sim$ reads as statistically distributed as. The symbol $\Re\{\cdot\}$ is the real-part operator.
@@ II System Model
In this scheme, each user’s message is split at the transmitter into a private and a common component, allowing receivers to partially decode interference while treating the remaining interference as noise. Specifically, the message $W_{k}$ (intended for user $k$) is divided into a private part $W_{\textrm{p}k}$ and a common part $W_{\textrm{c}k}$, for all $k\in\mathcal{K}$. These components are assumed to be statistically independent. The common parts $W_{\textrm{c}k}$ are aggregated into a single common message $W_{\rm c}$, which, along with the $K$ private messages $W_{\textrm{p}k}$, is independently encoded into streams $s_{\rm c}$ and $s_{k}$, respectively, where $\mathbb{E}\{|s_{k}|^{2}\}=1$. All streams are then linearly precoded prior to transmission. Hence, the transmit signal is written as:
@@
$y_{k}=\mathcal{L}_{k}\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{x}+n_{k},\quad k\in\mathcal{K},$ (2)
@@
$\mathbf{h}_{k}[m]=\sqrt{\epsilon^{2}}\,\mathbf{h}_{k}[m-1]+\sqrt{1-\epsilon^{2}}\,\mathbf{e}_{k}[m],$ (3)
@@
$\mathbf{P}_{k}=\mathbf{H}(\mathbf{H}^{\rm{H}}\mathbf{H})^{-1},$ (4)
where $\mathbf{H}=[\mathbf{h}_{1},...,\mathbf{h}_{K}]$ is a $N_{\rm{t}}\times K$ channel matrix, and $\mathbf{P}_{k}=[\mathbf{p}_{1},...,\mathbf{p}_{K}]$ is a $N_{\rm{t}}\times K$ precoder matrix. The common precoder is assumed to be a random beamformer independent of $\mathbf{h}_{k}[m-1]$, $\mathbf{e}_{k}[m]$, and $\mathbf{p}_{k}$, $\forall k\in\mathcal{K}$.
The instantaneous SNRs for the common and private streams at user $k$ are formulated as:
@@ III Main results
Without loss of generality and to simplify the notation, we will focus on the scenario of equal power allocation among users, i.e. $\alpha_{k}=\tau/K$, $\forall k\in\mathcal{K}$. Replacing this definition in (1), we obtain that the power weighting term of the private stream is $\tau$, since $\sum_{k\in\mathcal{K}}\alpha_{k}=\tau$ and the weight of the common part is $\alpha_{\rm c}=1-\tau$. Therefore, $\tau$ represents the total fraction of power allocated to private streams, with the case $\tau=1$ corresponding to the conventional SDMA scheme.
@@
${R}_{k}=\mathbb{E}\left\{\log_{2}\left(1+\frac{\mathcal{S}_{k}\tau}{K}X\right)\right\}-\mathbb{E}\left\{\log_{2}\left(1+\frac{\mathcal{S}_{k}\tau}{K}Z\right)\right\},$ (10)
@@
$X=\sum_{j\in\mathcal{K}}|\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2},$ (11)
$\textrm{Z}=\sum_{j\in\mathcal{K},j\neq k}|\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}.$ (12)
@@
$X=(1-\epsilon^{2})\sum_{j\in\mathcal{K}}|\mathbf{e}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}+\epsilon^{2}|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}+X_{\rm{CC}},$ (13)
@@
$X_{\rm{CC}}=2\sqrt{\epsilon^{2}(1-\epsilon^{2})}\Re(\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}\mathbf{e}_{k}^{\rm{H}}\mathbf{p}_{k})$ (14)
is a cross-correlation term. Due to the ZF precoder with uniform power allocation and the nature of the Rayleigh fading channel, the term $|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}$$\sim$ $\mathcal{G}(N_{\textrm{t}}-K+1,1)$}. Since the ZF precoder is independent of the Gaussian error, we have $|\mathbf{e}_{j}^{\rm{H}}[m]\mathbf{p}_{k}|^{2}\sim\mathcal{G}(1,1)$} [4]. Following a similar rationale, we rewrite (12) as:
$Z=\sum_{j\in\mathcal{K},j\neq k}|\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}\sim\mathcal{G}(K-1,1-\epsilon^{2}).$ (15)
@@
$D=\frac{(K(1-\epsilon^{2})+(N_{\rm t}-K+1)\epsilon^{2})^{2}}{\epsilon^{4}(N_{\rm t}+1)+(1-2\epsilon^{2})K+\mu},$ (16)
@@
$\Theta=\frac{\epsilon^{4}(N_{\rm t}+1)+(1-2\epsilon^{2})K+\mu}{K(1-\epsilon^{2})+(N_{\rm t}-K+1)\epsilon^{2}},$ (17)
where $\mu$ is:
$\mu=2\epsilon^{2}(1-\epsilon^{2})(N_{\rm t}-K+1).$ (18)
@@
Expressions (16) and (17) have been algebraically manipulated to match the structure of known ones [4, 6], facilitating comparison. Hence, the effect of neglecting the cross-correlation term in (14) is captured by parameter $\mu$ in (18).
@@
Interestingly, the mean of the Gamma distribution $\mathbb{E}\{X_{\rm{G}}\}=D\Theta$ does not depend on $\mu$. However, the variance $\mathrm{Var}\{X_{\rm{G}}\}=D\Theta^{2}$ does. Since $N_{\mathrm{t}}>K$, neglecting $\mu$ implies underestimating the true variance of $X$. This is akin to considering that the fading severity of the equivalent channel $X$ is lower than in practice.
@@ IV Numerical Results
In this section, we validate our analytical results by numerical performance evaluation against Monte Carlo (MC) simulations. First, we focus on providing evidence of the aspects discussed in Remarks 1 and 2. For this purpose, in Fig. 1 we represent the PDF of $X$ in (13), and provide comparisons with the Gamma approximations in Lemma 1 and [4, 6], referred [^1] Under uniform power allocation, both expressions are formally equivalent. to in the sequel as $X_{\rm{G}}$ and $X_{\rm{D}}$, respectively. The cases with $N=256$ $K=8$ users, CSIT staleness parameter $\epsilon=0.5$, and $N=32$, $K=8$ and $\epsilon=0.6$ are considered. As stated in Remark 2, the mean value is accurately captured by both approximations, being the first order statistics of the three distributions of interest well-aligned. However, we observe that the Gamma approximation in [4] largely underestimates the variance; this becomes more noticeable as the excess degrees of freedom captured by $(N_{\rm{t}}-K)$ grows. Conversely, the improved Gamma approximation in Lemma 1 closely matches the true distribution of $X$.
Figure 1: Empirical PDF of $X$ and approximated PDFs of $X_{\rm{G}}$ (Lemma 1) and $X_{\rm{D}}$ [4], for different values of $N_{\rm t}$, $K$ and $\epsilon$.
@@
where $M$ is the number of points on which the empirical PDF is estimated, ${\boldsymbol{\lambda}}=\left[D,\Theta\right]$ represents the parameters for the corresponding Gamma approximation, and $x_{k}$ correspond to the abscissa coordinates for such estimations. The MSE is evaluated in Fig. 2, as a function of the ratio $N_{\rm t}/K$, considering different numbers of users. We observe that the MSE provided by the approximation in [13] is consistently larger than that obtained by the approximation in Lemma 1. While the MSE of the former remains relatively constant as $N_{\rm t}/K$ increases, the MSE of the latter approximation decreases with $N_{\rm t}/K$, with a performance difference over two orders of magnitude. Hence, Lemma 1 provides a useful approximation in the operational regimes of practical MU-MIMO systems, i.e., $N_{\rm t}>>K$.
Figure 3: Ergodic sum-rate as a function of the power allocation factor $\tau$. The exact sum-rate obtained by (9), in solid blue line, is compared to the sum-rates obtained by the proposed Gamma approximation ($X_{\rm G}$) and the existing Gamma approximation ($X_{\rm D}$) in [4, 6]. The lower bound (LB) proposed in [4] is also evaluated when using ($X_{\rm G}$) and ($X_{\rm D}$) approximations. Parameter values are $N_{\rm t}=16$, $K=4$, with $\epsilon=\{0.3,0.5\}$
@@
There is one side effect of the variance underestimation of the approaches in [4, 6]: if one wishes to obtain lower bounds for the achievable rates, it is not possible to formally ensure that they always correspond to lower bounds if the starting point is actually an upper bound of the rate. Now, using the lower bounds for the achievable rates in [4, eq. (34)] with the improved values of $D$ and $\Theta$ in our approximation, we see that such bounds are looser than originally envisioned.

### v2

@@ Abstract
00footnotetext: This work has been submitted to the IEEE for publication. Copyright may be transferred without notice, after which this version may no longer be accessible00footnotetext: Manuscript received February 5, 2025; revised XXX. This work was supported in part by grant PID2023-149975OB-I00 (COSTUME) funded by MICIU/AEI/10.13039/501100011033 and FEDER/UE, in part by grant DGP_PRED_2024_0212 funded by Consejería de Universidad, Investigación e Innovación of Junta de Andalucía and the European Union, and in part by grant PMA1-2025-14-017 funded by Asociación Universitaria Iberoamericana de Postgrado. The review of this paper was coordinated by $XX$. (Corresponding author: M.C. Fernández Montefiore)00footnotetext: M.C. Fernández Montefiore and F.J. López-Martínez are with the Dept. Signal Theory, Networking and Communications, Research Centre for Information and Communication Technologies (CITIC-UGR), University of Granada, 18071, Granada, Spain. E-mail:{mcfm,fjlm}@ugr.es. M.C. Fernández Montefiore, G. González and F. Gregorio are with Instituto de Investigaciones en Ingeniería Eléctrica “Alfredo Desages” (IIIE), UNS-CONICET, and Departamento de Ingeniería Eléctrica y de Computadoras, Universidad Nacional del Sur (UNS), Argentina. E-mail: {ggonzalez,fernando.gregorio}@.uns.edu.ar00footnotetext: Digital Object Identifier 10.1109/XXX.2025.XXXXXXX
@@ I Introduction
In MU-MIMO systems, rate-splitting multiple access (RSMA) is proposed as an effective strategy to overcome the challenges posed by imperfect CSIT [6, 7]. In contrast to conventional multiple access techniques like space-division multiple access (SDMA), which treats interference as noise, and non-orthogonal multiple access (NOMA), leaning on interference decoding and cancellation, RSMA splits user messages into common and private parts, providing a more flexible approach [8] and a seamless integration with multiantenna operation. While private parts are dealt individually, the common message is decoded by all users. This allows RSMA to achieve a balance between interference suppression and decoding [9]. Furthermore, depending on the network topology and channel state condition, RSMA can adjust the power and message allocation based on interference levels, naturally transitioning between SDMA and NOMA [1, 10]. Such flexibility in managing interference makes RSMA more robust than conventional approaches, enabling a better mitigation of multi-user interference [11, 12] even in the presence of imperfect CSIT.
@@
Notation: $f_{(\cdot)}(\cdot)$ denotes a probability density function (PDF); $a$ is a scalar, $\mathbf{a}$ is a vector, and $\mathbf{A}$ is a matrix. Conjugate transpose of $\mathbf{A}$ is denoted by $\mathbf{A}^{\rm{H}}$. $\overline{a}$ denotes the complex conjugate of $a$. The expectation operator is denoted as $\mathbb{E}\left\{\cdot\right\}$, and $\mathrm{Var}\{\cdot\}$ denotes variance. $\mathcal{C}\mathcal{N}(\mu,\sigma^{2})$ represents the complex Gaussian distribution with parameters $\mu$ and $\sigma^{2}$; $\mathcal{G}(D,\Theta)$ represents the Gamma distribution with shape and scale parameters $D$ and $\Theta$, respectively. The symbol $\sim$ reads as statistically distributed as. The symbol $\Re\{\cdot\}$ is the real-part operator.
@@ II System Model
In this scheme, each user’s message is split at the transmitter into a private and a common component, allowing receivers to partially decode interference while treating the remaining interference as noise. Specifically, the message $W_{k}$ (intended for user $k$) is divided into a private part $W_{\textrm{p}k}$ and a common part $W_{\textrm{c}k}$, for all $k\in\mathcal{K}$. The common parts $W_{\textrm{c}k}$ are aggregated into a single common message $W_{\rm c}$. Then, $W_{\rm c}$ along with the $K$ private messages $W_{\textrm{p}k}$ are independently encoded into streams $s_{\rm c}$ and $s_{k}$, respectively, where $\mathbb{E}\{|s_{k}|^{2}\}=1$. All streams are then linearly precoded prior to transmission. Hence, the transmit signal is written as:
@@
$y_{k}=\sqrt{\mathcal{L}_{k}}\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{x}+n_{k},\quad k\in\mathcal{K},$ (2)
@@
$\mathbf{h}_{k}[m]=\sqrt{\epsilon^{2}}\,\mathbf{h}_{k}[m-1]+\sqrt{1-\epsilon^{2}}\,\mathbf{e}_{k}[m].$ (3)
@@
$\mathbf{P}=\mathbf{H}(\mathbf{H}^{\rm{H}}\mathbf{H})^{-1},$ (4)
where $\mathbf{H}=[\mathbf{h}_{1},...,\mathbf{h}_{K}]$ is a $N_{\rm{t}}\times K$ channel matrix, and $\mathbf{P}=[\mathbf{p}_{1},...,\mathbf{p}_{K}]$ is a $N_{\rm{t}}\times K$ precoder matrix. The common precoder is assumed to be a random beamformer independent of $\mathbf{h}_{k}[m-1]$, $\mathbf{e}_{k}[m]$, and $\mathbf{p}_{k}$, $\forall k\in\mathcal{K}$.
The instantaneous SINRs for the common and private streams at user $k$ are formulated as:
@@ III Main results
${R}_{k}=\mathbb{E}\left\{\log_{2}\left(1+\mathcal{S}_{k}X\right)\right\}-\mathbb{E}\left\{\log_{2}\left(1+\mathcal{S}_{k}Z\right)\right\},$ (10)
@@
$X=\sum_{j\in\mathcal{K}}\alpha_{j}|\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2},$ (11)
$\textrm{Z}=\sum_{j\in\mathcal{K},j\neq k}\alpha_{j}|\mathbf{h}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}.$ (12)
@@
$\hskip-11.66563ptX=(1-\epsilon^{2})\sum_{j\in\mathcal{K}}\alpha_{j}|\mathbf{e}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}+\epsilon^{2}\alpha_{k}|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}+X_{\rm{CC}},$ (13)
@@
$X_{\rm{CC}}=2\sqrt{\epsilon^{2}(1-\epsilon^{2})}\alpha_{k}\Re(\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}\mathbf{e}_{k}^{\rm{H}}[m]\mathbf{p}_{k})$ (14)
is the cross-correlation term between the effective channel gain and the projection of the channel error into the direction of the user-$k$. Due to the ZF precoder with uniform power allocation and the nature of the Rayleigh fading channel, the term $|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}$$\sim$ $\mathcal{G}(N_{\textrm{t}}-K+1,1)$}. Since the ZF precoder is independent of the Gaussian error, we have $|\mathbf{e}_{j}^{\rm{H}}[m]\mathbf{p}_{k}|^{2}\sim\mathcal{G}(1,1)$} [4]. Following a similar rationale, we rewrite (12) as:
$Z\sim\mathcal{G}\left(\frac{(1-\alpha_{c}-\alpha_{k})^{2}}{\sum_{j\in\mathcal{K},j\neq k}\alpha_{j}^{2}},\frac{(1-\epsilon^{2})\sum_{j\in\mathcal{K},j\neq k}\alpha_{j}^{2}}{1-\alpha_{c}-\alpha_{k}}\right).$ (15)
@@
$D=\frac{((N_{\rm t}-K+1)\epsilon^{2}\alpha_{k}+(1-\alpha_{c})(1-\epsilon^{2}))^{2}}{(N_{\rm t}-K+1)\epsilon^{4}\alpha_{k}^{2}+\sum_{j\in\mathcal{K}}\alpha_{j}^{2}(1-\epsilon^{2})^{2}+\mu_{k}},$ (16)
@@
$\Theta=\frac{(N_{\rm t}-K+1)\epsilon^{4}\alpha_{k}^{2}+\sum_{j\in\mathcal{K}}\alpha_{j}^{2}(1-\epsilon^{2})^{2}+\mu_{k}}{(N_{\rm t}-K+1)\epsilon^{2}\alpha_{k}+(1-\alpha_{c})(1-\epsilon^{2})}$ (17)
where $\mu_{k}$ is:
$\mu_{k}=\mathbb{E}\left\{X_{\rm{CC}}^{2}\right\}=2\epsilon^{2}\alpha_{k}^{2}(1-\epsilon^{2})(N_{\rm t}-K+1).$ (18)
@@
The expressions (16) and (17) have been algebraically manipulated to match the structure of known ones [4, 8], facilitating comparison. Hence, the effect of neglecting the cross-correlation term in (14) is captured by the parameter $\mu_{k}$ in (18), which is the power of $X_{\rm{CC}}$.
@@
Interestingly, the mean of the Gamma distribution $\mathbb{E}\{X_{\rm{G}}\}=D\Theta$ does not depend on $\mu_{k}$. However, the variance $\mathrm{Var}\{X_{\rm{G}}\}=D\Theta^{2}$ does. Since $N_{\mathrm{t}}>K$, neglecting $\mu_{k}$ implies underestimating the true variance of $X$. This is akin to considering that the fading severity of the equivalent channel $X$ is lower than in practice.
@@ IV Numerical Results
In this section, we validate our analytical results by numerical performance evaluation against Monte Carlo (MC) simulations. First, we focus on providing evidence of the aspects discussed in Remarks 1 and 2. For this purpose, in Fig. 1 we represent the PDF of $X$ in (13), and provide comparisons with the Gamma approximations in Lemma 1 and [4, 8] [^1] Under uniform power allocation, both expressions are formally equivalent., referred to in the sequel as $X_{\rm{G}}$ and $X_{\rm{D}}$, respectively. The cases with $N_{\rm t}=256$, $K=8$ users, CSIT staleness parameter $\epsilon=0.5$, and $N_{\rm t}=32$, $K=8$ and $\epsilon=0.6$ are considered; for simplicity, we use $\alpha_{k}=1/K$ so that $\mu_{k}=\mu,\forall k$. As stated in Remark 2, the mean value is accurately captured by both approximations, being the first order statistics of the three distributions of interest well-aligned. However, we observe that the Gamma approximations in [4, 8] largely underestimate the variance; this becomes more noticeable as the excess degrees of freedom captured by $(N_{\rm{t}}-K)$ grows. Conversely, the improved Gamma approximation in Lemma 1 closely matches the true distribution of $X$.
Figure 1: Empirical PDF of $X$ and approximated PDFs of $X_{\rm{G}}$ (Lemma 1) and $X_{\rm{D}}$ [4, 8], for different values of $N_{\rm t}$, $K$, and $\epsilon$.
@@
where $M$ is the number of points on which the empirical PDF is estimated, ${\boldsymbol{\lambda}}=\left[D,\Theta\right]$ represents the parameters for the corresponding Gamma approximation, and $x_{k}$ corresponds to the abscissa coordinates for such estimations. The MSE is evaluated in Fig. 2, as a function of the ratio $N_{\rm t}/K$, considering different numbers of users. We observe that the MSE provided by the approximations in [4, 8] is consistently larger than that obtained by the approximation in Lemma 1. While the MSE of the former remains relatively constant as $N_{\rm t}/K$ increases, the MSE of the latter approximation decreases with $N_{\rm t}/K$, with a performance difference over two orders of magnitude. Hence, Lemma 1 provides a useful approximation in the operational regimes of practical MU-MIMO systems, i.e., $N_{\rm t}>>K$.
Figure 3: Ergodic sum-rate as a function of the power allocation factor $1-\alpha_{c}$. The exact sum-rate obtained by (9), in solid blue line, is compared to the sum-rates obtained by the proposed Gamma approximation $X_{\rm G}$ and the existing Gamma approximation $X_{\rm D}$ in [4, 8]. The lower bound (LB) proposed in [4] is also evaluated when using $X_{\rm G}$ and $X_{\rm D}$ approximations. Parameter values are $N_{\rm t}=16$, $K=4$, with $\epsilon=\{0.3,0.5\}$
@@
There is one side effect of the variance underestimation of the approaches in [4, 8]: if one wishes to obtain lower bounds for the achievable rates, it is not possible to formally ensure that they always correspond to lower bounds if the starting point is actually an upper bound of the rate. Now, using the lower bound definitions for the achievable rates in [4, eq. (34)] with the improved values of $D$ and $\Theta$ in our approximation, we see that such bounds are looser than originally envisioned.